Q = L2 K. Calculate the elasticity of substitution


1
Expert's answer
2022-02-04T07:48:17-0500

Q = L2K

Elasticity of substitution ("\alpha") = "\frac{percentage change in K/L}{percentage change in MRTS}"


% change in K/L: = "{dIn(K\/L)}"

MRTS: ="\frac{MPL}{MPK}"


"MPL=\frac {\delta Q}{\delta L}=LK"


 "MPK=\frac{\delta Q}{\delta K }=L^{2}"


"MRTS=\frac{LK}{L^{2}}=\frac{K}{L}"


"\alpha" = "\frac{dIn(K/L)}{dIn(K/L)}"


"\alpha" = 1